OpenStudy (anonymous):

Choose the correct product of (2x − 4)2

OpenStudy (zale101):

(2x - 4)^2 is (2x - 4)(2x - 4) = 4x^2 - 8x - 8x + 16 =

OpenStudy (zale101):

use this formula (a - b)^2 = a^2 - 2*a*b + b^2

OpenStudy (mathstudent55):

Can you do FOIL? \( (2x - 4)^2 = (2x - 4)(2x - 4) \)

OpenStudy (anonymous):

4x^2-16x+16??

OpenStudy (zale101):

yes